A student throws a set of keys vertically upward to her sorority sister, who is in a window 3.70 m above. The keys are caught 1.30 s later by the sisters outstretched hand (a) With what initial velocity were the keys thrown? Your response differs from the correct answer by more than 10%. Double check your calculations. m/s upward (b) What was the velocity of the keys just before they were caught? m/s downward
Answer #1

Suppose initial velocity is U and final velocity is V

Now we know from kinematic equations that

V = U + a*t

H = U*t + 0.5*a*t^2

Given that

H = height between both student = 3.70 m

t = time taken = 1.30 sec

a = acceleration = -g = -9.81 m/sec

So using 2nd equation

3.70 = U*1.30 - 0.5*9.81*1.30^2

U = [3.70 + 0.5*9.81*1.3^2]/1.3

U = 9.22 m/sec

Part B

V = U + a*t

V = 9.22 - 9.81*1.30

V = -3.53 m/sec

|V| = 3.53 m/sec

-ve sign means 2nd student catches key when they are coming back down.

Please Upvote.

Comment below if you have any doubt.